### v2

# arXiv:2602.21157v2[cs.RO] 27 Feb 2026

Abstract
Vision–Language–Action (VLA) models have shown strong performance in robotic manipulation, but often struggle in long-horizon or out-of-distribution scenarios due to the lack of explicit mechanisms for multimodal reasoning and anticipating how the world will evolve under action. Recent works introduce textual chain-of-thought or visual subgoal prediction within VLA models to reason, but still fail to offer a unified human-like reasoning framework for joint textual reasoning, visual foresight, and action prediction. To this end, we propose Halo, a unified VLA model that enables embodied multimodal chain-of-thought (EM-CoT) reasoning through a sequential process of textual task reasoning, visual subgoal prediction for fine-grained guidance, and EM-CoT-augmented action prediction. We instantiate Halo with a Mixture-of-Transformers (MoT) architecture that decouples semantic reasoning, visual foresight, and action prediction into specialized experts while allowing seamless cross-expert collaboration. To enable Halo learning at scale, we introduce an automated pipeline to synthesize EM-CoT training data along with a carefully crafted training recipe. Extensive experiments demonstrate that: (1) Halo achieves superior performance in both simulated and real-world environments, surpassing baseline policy $\pi_{0}$ by 34.1% on RoboTwin benchmark; (2) all proposed components of the training recipe and EM-CoT design help improve task success rate; and (3) Halo exhibits strong generalization capabilities under aggressive unseen environmental randomization with our proposed EM-CoT reasoning.
[…]
1 Introduction
Vision-Language-Action (VLA) models (Black et al., 2024; NVIDIA et al., 2025) have shown a compelling path toward general-purpose robotic manipulation. However, most existing VLAs map perceptual inputs directly to motor commands, lacking explicit mechanisms for reasoning about task structure or anticipating how the environment will evolve under motion and contact. This limitation becomes particularly pronounced in long-horizon or out-of-distribution scenarios—such as novel layouts, unfamiliar objects, or contact-rich interactions—where successful execution depends more on deliberation and foresight than on reactive pattern matching.
Recent work has sought to address this limitation by introducing intermediate reasoning processes like human. Inspired by the success of Chain-of-Thought reasoning (Wei et al., 2023; Chen et al., 2025a, b) in large language models, Zawalski et al. (2025) train VLAs to perform multi-step reasoning prior to action prediction, enabling task decomposition but without grounding such reasoning in explicit visual state evolution. In contrast, Zhao et al. (2025) autoregressively generate subgoal images as visual intermediate steps, but lack semantic reasoning capabilities. Achieving human-like textual reasoning and visual imagination capability within a unified framework remains challenging, as it demands strong multimodal generation while preserving reasoning capacity. Existing approaches (Zhao et al., 2025; Gu et al., 2025) often tightly couple visual generation with VLMs via discrete visual token prediction, which may hinder rich understanding capabilities of VLMs. Consequently, a unified architecture that jointly supports multimodal reasoning, visual generation, and action prediction remains an open problem.
Figure 1: Halo first performs textual reasoning and task planning, then predicts visual subgoals for fine-grained guidance, and finally generates actions conditioned on EM-CoT. This process is implemented using a Mixture-of-Transformers (MoT) architecture that integrates a multimodal understanding expert, a visual generation expert, and an action prediction expert together through shared self-attention.
To address this, we propose Halo, a unified VLA model that enables embodied multimodal chain-of-thought (EM-CoT) reasoning. As illustrated in Figure 1, Halo first performs textual reasoning and task planning, then predicts subgoal images to provide fine-grained visual guidance, and finally uses the resulting EM-CoT as context for action prediction. This design is specifically engineered to emulate the human cognitive process of deliberate reasoning, mental imagination, and subsequent execution, rendering Halo particularly well-suited for tackling complex, long-horizon tasks that require intricate sequential reasoning and sustained contextual understanding over extended periods.
Halo incorporates several key innovations to realize this vision. First, to natively support embodied multimodal reasoning, Halo adopts a Mixture-of-Transformers (MoT) architecture (Liang et al., 2024) that decouples textual reasoning, visual foresight, and action prediction into three specialized experts. These experts collaborate seamlessly through shared self-attention, while preserving each expert’s natural generative workflow—autoregressive text generation for reasoning and diffusion-based prediction for visual states and actions—thereby avoiding potential conflicts that arise when heterogeneous capabilities are forced into a single monolithic model. Second, we introduce an automatic EM-CoT data synthesis pipeline built upon action primitives and VLM annotator. The pipeline performs task planning first and then decomposes long-horizon robot trajectories into a sequence of sub-tasks, each paired with a corresponding subgoal image and explicit reasoning process. This enables automatic and scalable embodied multimodal reasoning supervision for our unified VLA model. Third, we propose a carefully designed training recipe that combines broad generalization with embodied reasoning specialization. (1) In a versatile pre-training stage, Halo is trained on diverse carefully crafted data sources, including general VQA datasets for robust multimodal understanding, egocentric and robotic video datasets for learning contact-rich physical dynamics, and heterogeneous robot trajectory datasets for acquiring foundational manipulation skills. (2) In a subsequent EM-CoT–augmented fine-tuning stage, the model is co-trained on EM-CoT and VQA data to inject EM-CoT reasoning capability while preserving its general world knowledge.
We conduct extensive experiments in both the RoboTwin 2.0 simulator (Chen et al., 2025c) and real-world environments to validate the effectiveness of Halo. The results demonstrate that Halo achieves consistently stronger overall performance across both simulated and real-world settings, establishing a new state-of-the-art with an average success rate of 80.5% on RoboTwin, surpassing $\pi_{0}$ by 34%. In addition, we perform comprehensive ablation studies to verify: (1) the effectiveness of the proposed EM-CoT, showing that combining textual reasoning with visual foresight contribute significantly to performance; and (2) the importance of each category of pre-training tasks, all of which provide measurable benefits to downstream performance. We further demonstrate the strong generalization capability of Halo in both simulation and real robots, and qualitatively show that it can generate reasonable EM-CoT to guide action prediction in unseen scenarios. Taken together, these results highlight Halo as a scalable and generalizable paradigm for human-like VLA reasoning.
2 Related Work
Vision-Language-Action Models.
Predominant VLA models are built on top of Vision-Language Models (VLMs) pre-trained on large-scale web data, inheriting rich multimodal knowledge about the world. Representative architectures such as OpenVLA (O’Neill et al., 2024) and FAST (Pertsch et al., 2025) directly extend pre-trained VLMs by introducing discrete action tokens and predicting actions autoregressively to solve embodied tasks. The $\pi_{0}$ series (Black et al., 2024; Intelligence et al., 2025) further introduce a separate action expert along with the pre-trained VLM for flow-based action prediction, achieving better performance and efficiency. Beyond equipping VLM with only action prediction capability, unified VLA architectures (Bi et al., 2025; Wang et al., 2025; Cen et al., 2025; Lv et al., 2025) have been proposed to enhance dynamic awareness by jointly modeling textual understanding, visual generation, and action prediction. These methods incorporate image or video generation components (Zhu et al., 2025, 2026) to predict future environmental states, using either unified (Cen et al., 2025; Wang et al., 2025) or mixed (Bi et al., 2025; Lv et al., 2025) architectures that are trained together with action prediction to capture underlying environment dynamics. However, these methods lack explicit textual or visual reasoning capability, limiting their performance in complex and dynamic environments.
Textual and Visual Reasoning for Robotic Intelligence.
Reasoning is a crucial step towards robotic intelligence (LeCun, 2022). Motivated by the success of Chain-of-Thought reasoning in language models, recent work extends explicit reasoning mechanisms to embodied decision making. MoTVLA (Bi et al., 2025) proposes a textual reasoning framework that performs planning and reasoning prior to action execution. UP-VLA (Zhang et al., 2025), CoTVLA (Zhao et al., 2025) and InternVLA-A1 (Cai et al., 2026) further introduce visual reasoning by generating goal images and videos to guide downstream action prediction. More recently, ManualVLA (Gu et al., 2025) tries to combine visual and textual reasoning through a shared planning expert that produces both textual instructions and visual subgoals. However, this design couples image and text generation within a single autoregressive model, which may degrade the reasoning capability of the underlying VLM. In contrast, our proposed embodied multimodal chain-of-thought (EM-CoT) employs a Mixture-of-Transformers (MoT) architecture for joint textual reasoning, visual generation, and action prediction. Our decoupled design enables more effective reasoning across visual and textual modalities before executing actions.
3 Method
In this section, we delineate the core design principles of Halo, focusing on its unified architecture, the embodied multimodal chain-of-thought (EM-CoT) data pipeline, and the associated training recipe. Collectively, these components empower the model to engage in textual reasoning, visual foresight, and grounded action prediction.
3.1 Problem Formulation
We formulate robotic manipulation as a sequential decision-making problem. Let $\tau=\{(\mathbf{o}_{t},\mathbf{l},\mathbf{a}_{t})\}_{t=1}^{T}$ denote a trajectory, comprising visual observations $\mathbf{o}_{t}\in\mathcal{O}$, language instructions $\mathbf{l}\in\mathcal{L}$, and continuous actions $\mathbf{a}_{t}\in\mathcal{A}$. Traditional VLA models typically learn a monolithic policy $\pi_{\theta}(\mathbf{a}_{t:t+m}\mid\mathbf{l},\mathbf{o}_{t-k:t})$ that directly maps history observations and instructions to action chunks. Such purely reactive policies often suffer from performance degradation when facing long-horizon or complex manipulation tasks due to a lack of intermediate reasoning.
In this case, we aim to integrate explicit EM-CoT as a robust conditioning mechanism in Halo and make its generation process align with typical human cognitive pathway of “thinking-imagination-execution”. To achieve our goal, we decompose the generation process of Halo into three distinct phases: textual reasoning, visual foresight, and action prediction. Then the problem is transformed into learning a unified policy $\pi_{\theta}$ that jointly models the following mappings:
$\mathbf{r}$ $\sim P_{\theta}(\cdot\mid\mathbf{l},\mathbf{o}_{t-k:t})$ (1)
$\hat{\mathbf{o}}_{t+h}$ $\sim P_{\theta}(\cdot\mid\mathbf{l},\mathbf{o}_{t-k:t},\mathbf{r})$ (2)
$\mathbf{a}_{t:t+m}$ $\sim\pi_{\theta}(\cdot\mid\mathbf{l},\mathbf{o}_{t-k:t},\mathbf{r},\hat{\mathbf{o}}_{t+h})$ (3)
where $\mathbf{r}$ represents generated textual chain-of-thought and $\hat{\mathbf{o}}_{t+h}$ denotes predicted visual subgoal. Under this formulation, the action chunk $\mathbf{a}_{t:t+m}$ is not a simple reaction, but the result of a deliberate reasoning trace and physical world modeling, ensuring that every control command is strategically planned and visually grounded.
3.2 Unified Architecture
We now present the unified architecture of Halo, as illustrated in Figure 1. Drawing inspiration from the capability of BAGEL (Deng et al., 2025) to harmonize multimodal tasks, we adopt a Mixture-of-Transformers (MoT) (Liang et al., 2024) architecture comprising three specialized experts: Multimodal Understanding, Visual Generation, and Action Prediction. These three experts specialize in distinct modalities and are responsible for textual reasoning, visual generation, and action prediction, respectively. While maintaining independent parameter sets, they interact through a shared self-attention mechanism to enable rich cross-modal interactions. With such an architecture, Halo operates as a unified VLA model that simultaneously possesses the capability to generate detailed textual reasoning and to produce explanatory subgoal images.
The switching mechanism between modalities is explicitly controlled via special tokens. By default, the model operates as an auto-regressive planner; however, the generation of specific tokens (e.g., $\langle\text{visual\_start}\rangle$ or $\langle\text{action\_start}\rangle$) triggers the routing of hidden states to the visual generation expert or the action prediction expert, respectively. In order to manage the information flow among experts, we implement a carefully structured masking strategy within the shared self-attention. As illustrated in Figure 4, a causal mask is applied to textual tokens to enforce autoregressive generation. Conversely, visual tokens employ bidirectional attention within each frame to capture global spatial dependencies, while maintaining causal masking for interactions across frames or modalities. Crucially, to prevent information leakage, noise tokens are restricted from attending to their corresponding ground-truth targets. Furthermore, all other tokens are masked from attending to noise tokens, ensuring they only aggregate information from valid non-noise contexts.
To facilitate training stability, we initialize each expert using the Qwen2.5-1.5B LLM (Yang et al., 2025), yielding a cumulative parameter count of approximately 4.5B. To project heterogeneous data into a unified representation space, we employ modality-specific encoders. For text, we adopt the standard Qwen2.5 tokenizer (Bai et al., 2023), expanding the vocabulary with special control tokens—specifically $\langle\text{think\_start}\rangle$, $\langle\text{vision\_start}\rangle$, and $\langle\text{action\_start}\rangle$, among others—to strictly demarcate the textual reasoning, visual generation, and acting phases, thereby enforcing a structured and interpretable workflow. In the vision domain, we decouple the encoding mechanisms for semantic understanding and visual generation to address their distinct functional requirements. For the understanding task, we employ a ViT (Dosovitskiy, 2020) initialized with SigLIP2 (Tschannen et al., 2025) to leverage robust semantic features. To accommodate diverse real-world inputs without geometric distortion, this encoder is further enhanced with NaViT (Dehghani et al., 2023), enabling the processing of images in their native aspect ratios. For the visual generation task, we utilize a pre-trained VAE (Kingma & Welling, 2013; Labs et al., 2025) optimized for high-fidelity image compression and reconstruction. Finally, for the action modality, given its low-dimensional and physically structured nature, a linear projection suffices to map actions into the model’s hidden dimension while preserving intrinsic structural information. See Appendix A for further architectural details.
Figure 2: Overview of EM-CoT Data Synthesis Pipeline. The pipeline converts raw robotic trajectories into EM-CoT data in three phases: (1) action primitives are extracted from robot proprioception via rule-based matching; (2) a VLM acts as an annotator to generate task plans, decompose trajectories into subtasks, and align each subtask with explicit textual reasoning; and (3) the terminal frame of each subtask is selected as a visual subgoal image, producing structured embodied multimodal chain-of-thought supervision.
3.3 EM-CoT Data Pipeline
To support the generation of intermediate textual reasoning trace $\mathbf{r}$ and visual subgoal $\hat{\mathbf{o}}_{t+h}$, we construct an automatic pipeline to synthesize EM-CoT data from raw robot trajectories at scale. As shown in Figure 2, the pipeline operates in three phases. Firstly, we translate continuous low-level actions into high-level motion primitives via rule-based matching following Belkhale et al. (2024). Secondly, we utilize a large-scale Vision-Language Model (Qwen3-VL (Bai et al., 2025)) to augment the trajectory with dense textual reasoning, including task narratives and subtask decomposition. Finally, the terminal frame of each subtask is designated as its corresponding visual subgoal, providing a sparse supervision signal that effectively lowers the learning difficulty. This process yields a high-quality dataset $\mathcal{D}_{\text{ft}}$ where every trajectory is augmented with aligned textual reasoning $\mathbf{r}$ and visual goals $\hat{\mathbf{o}}$, enabling the supervision of the intermediate steps defined in Eq. (1)-(2). See Appendix B for more details on this pipeline, including the pseudo code and tailored prompt used in each phase.
Figure 3: Attention Masking Strategy for EM-CoT. (1) Spatial and semantic tokens utilize bidirectional attention within frames. (2) Noise tokens attend bidirectionally to each other. (3) Cross-modality interactions and text generation follow causal constraints. (4) Non-noise tokens are masked from noise tokens to prevent leakage.
Figure 4: Overview of dataset recipe. Halo training involves two stages: Stage 1 pre-trains on general VQA, visual generation, and action prediction to build foundation skills; Stage 2 introduces EM-CoT–augmented fine-tuning to inject multi-step reasoning and foresight capabilities.
3.4 Training Recipe
To evolve Halo into a capable deliberative unified VLA, we employ a two-stage training paradigm: (1) Versatile Pre-training, which establishes a robust generalist foundation, and (2) EM-CoT-Augmented Fine-tuning, which elicits structured multimodal reasoning capabilities.
Stage 1: Versatile Pre-training.
The primary objective of this phase is to unify multimodal understanding, physical dynamics prediction, and foundational manipulation skills into a single architecture. To this end, we curate a massive, heterogeneous dataset spanning three distinct domains—Visual Question Answering (VQA), Visual Generation (VG), and Action Prediction (AP)—as illustrated in Fig. 4. This diversity ensures the model develops a generalized representation capable of supporting complex downstream reasoning.
  • VQA (Mutilmodal understanding): We use LLaVA-NeXT-779k (Liu et al., 2024a) to provide high-quality multimodal conversational priors. This task aligns linguistic instructions with visual contexts via a cross-entropy loss, denoted as $\mathcal{L}_{\text{CE}}$.
  • VG (Visual Generation): To instill physical common sense, we integrate robotic trajectories from OXE (O’Neill et al., 2024) with ego-centric manipulation videos from SSv2 (Goyal et al., 2017). We formulate visual foresight as a future frame prediction task, $(l,I_{t-k:t})\to I_{t+h}$, trained using a flow-matching MSE loss, $\mathcal{L}_{\text{MSE}}$. Crucially, we employ a dual-path visual pathway that integrates complementary semantic and spatial representations: a ViT branch first captures high-level semantic context, while a VAE branch provides spatially grounded latent features that are compact and generation-friendly. By combining these two pathways, the model jointly exploits semantic dynamics and spatial structure for prediction.
  • AP (Imitation Learning): We repurpose the OXE dataset for direct action prediction. The model learns to predict the next action chunk $a_{t}$ given the history, supervised by an $L_{1}$ flow-matching loss, $\mathcal{L}_{\text{L1}}$.
To address the varying optimization difficulties across these tasks, we balance the pre-training objective by assigning higher weights to manipulation intensive tasks. The total pre-training loss is defined as:
$\mathcal{L}_{\text{pt}}=0.25\mathcal{L}_{\text{CE}}+0.5\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{L1}}.$ (4)
Table 1: Quantitative results in simulation on RoboTwin 2.0. Each task is evaluated 100 times under the Easy (Clean) and Hard (Domain-randomized) settings, and the average success rate across 50 tasks is reported. The complete results are provided in Appendix D.
Task $\pi_{0}$ RDT-1B Diffusion Policy Halo-w/o EM-CoT Halo
Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard
Adjust Bottle 90 56 81 75 97 0 99 2 97 9
Beat Block Hammer 43 21 77 37 42 0 95 3 96 11
Blocks Ranking RGB 19 5 3 0 0 0 85 3 94 7
Blocks Ranking Size 7 1 0 0 1 0 37 0 58 2
Click Alarmclock 63 11 61 12 61 5 85 10 83 14
Click Bell 44 3 80 9 54 0 100 5 100 10
Dump Bin Bigbin 83 24 64 32 49 0 88 30 93 28
Grab Roller 96 80 74 43 98 0 97 49 95 57
$\cdots$
Shake Bottle Horizontally 99 51 84 51 59 18 100 63 100 66
Shake Bottle 97 60 74 45 65 8 100 70 98 73
Stack Blocks Three 17 0 2 0 0 0 90 32 96 37
Stack Blocks Two 42 1 21 2 7 0 99 55 100 60
Stack Bowls Three 66 24 51 17 63 0 83 20 92 25
Stack Bowls Two 91 41 76 30 61 0 96 46 98 49
Stamp Seal 3 4 1 0 2 0 49 16 60 21
Turn Switch 27 23 35 15 36 1 58 24 65 27
Average 46.4 16.3 34.5 13.7 28.0 0.6 75.3 21.2 80.5 26.4
Stage 2: EM-CoT-Augmented Fine-tuning.
Building upon the pre-trained foundation, we transition Halo from a reactive policy to a deliberative one by fine-tuning it on our EM-CoT dataset, $\mathcal{D}_{\text{ft}}$ (see Sec. 3.3). This stage augments standard robotic trajectories with explicit textual reasoning traces $\mathbf{r}$ and visual subgoals $\hat{\mathbf{o}}$.
Adhering to the sequential structure defined in Eq. (1)–(3), the model is trained to orchestrate a complete “thought-foresight-action” chain. The fine-tuning objective minimizes the joint loss:
$\mathcal{L}_{\text{ft}}=\mathcal{L}_{\mathbf{r}}+\mathcal{L}_{\hat{\mathbf{o}}}+\mathcal{L}_{\mathbf{a}},$ (5)
where $\mathcal{L}_{\mathbf{r}}$, $\mathcal{L}_{\hat{\mathbf{o}}}$, and $\mathcal{L}_{\mathbf{a}}$ represent the losses for textual reasoning, visual subgoal generation, and action prediction, respectively. To mitigate catastrophic forgetting of general multimodal capabilities during this specialization, we incorporate general VQA data into the fine-tuning corpus (Cheang et al., 2025). This process results in a model that effectively bridges high-level semantic intent with precise, visually grounded motor control for long-horizon manipulation. Please refer to Appendix C for implementation details.
4 Experiments
We conduct extensive experiments to evaluate the effectiveness of Halo equipped with EM-CoT. Our study focuses on: (i) whether the proposed unified VLA architecture with EM-CoT improves overall performance and generalization; (ii) whether Halo can generate informative EM-CoT reasoning, particularly under novel settings; (iii) whether multimodal reasoning outperforms using only textual reasoning or visual subgoal prediction, and how the proposed training recipe contributes to performance; and (iv) whether Halo demonstrates superior performance in real-world robotic settings.
4.1 Experiment Settings
During pre-training, following (Deng et al., 2025), training samples are concatenated to a maximum sequence length of 27k tokens, and Flex Attention (Dong et al., 2024) is adopted to accelerate training; Halo is trained for 90k steps at this stage. For fine-tuning, we use raw trajectories collected from RoboTwin 2.0 (Chen et al., 2025c) and real-world. The simulation dataset contains 2,500 expert demonstrations (50 per task) collected in clean environments, while the real-world dataset consists of 320 demonstrations (80 per task). Using these datasets, Halo is fine-tuned for 110k steps in simulation and 80k steps in real-world experiments. During inference, task instructions together with $c=3$ observation frames are provided as input. We set the action chunk length to $K=16$ for simulation and $K=50$ for real-world experiments.
4.2 Simulation Results
We conduct simulation experiments on RoboTwin 2.0 (Chen et al., 2025c), a comprehensive benchmark comprising 50 challenging manipulation tasks. We compare Halo with several competitive baselines, including $\pi_{0}$ (Black et al., 2024), RDT (Liu et al., 2024b), and Diffusion Policy (Chi et al., 2025). Baseline results are taken directly from the official RoboTwin 2.0 leaderboard. [^1] https://robotwin-platform.github.io/leaderboard To isolate the contribution of EM-CoT, we further include a variant of our method without EM-CoT, denoted as Halo-w/o EM-CoT.
Table 1 shows the performance of Halo and all baselines on RoboTwin 2.0. It can be observed that Halo consistently outperforms all competitive baselines across both Easy and Hard settings. The average success rate of Halo reaches 80.46% on Easy tasks and 26.44% on Hard tasks, surpassing baseline policy $\pi_{0}$ by 34.1% and 10.1%, respectively. Particularly, the consistent huge relative performance gap (i.e., 73.5% and 62.0%) between Halo and $\pi_{0}$ especially on Hard tasks indicates that Halo can also handle out-of-distribution (OOD) challenges more robustly. This stems from the fact that Halo not only possesses robust foundational capabilities but also features EM-CoT capabilities meticulously tailored for intricate, complex scenarios. Besides, even without the explicit reasoning chain, Halo-w/o EM-CoT already surpasses the strongest baseline ($\pi_{0}$) by a substantial margin of +28.92+28.92 points on average for Easy tasks, demonstrating the powerful general foundation established by our versatile pre-training.
In contrast, traditional reactive policies such as Diffusion Policy struggle significantly in randomized environments, with success rates dropping to near-zero (0.6%) in Hard settings. While $\pi_{0}$ and RDT-1B exhibit some degree of robustness, they plateau at significantly lower performance levels, whereas Halo continues to improve as it incorporates more structured reasoning. On fine-grained tasks such as “Blocks Ranking Size” and “Stamp Seal,” Halo achieves a multi-fold increase in success rates compared to baselines, demonstrating its ability to manage intricate object interactions. These results underscore the effectiveness and robustness of Halo for complex robotic manipulation.
Table 2: Ablation studies on training recipe and EM-CoT. Panel A: training recipe ablation (V/T/A denote visual generation, textual VQA, and action prediction data. Note that in the Full setting, we pre-train the model with V&T&A datasets and then fine-tune it on the official RoboTwin 2.0 dataset). Panel B: EM-CoT ablation (V = visual subgoal images, T = textual reasoning).
Panel A: Training Recipe Panel B: EM-CoT
Setting Easy Hard Setting Easy Hard
Full 75.3 21.2 Halo 80.5 26.4
w/o V 58.2 10.5 w/o T 77.8 18.3
w/o V+T 42.9 3.9 w/o V 76.1 22.5
w/o V+T+A 32.4 0.0 w/o V & T 75.3 21.2
4.3 Ablation Study
We perform ablation studies to validate the effectiveness of Halo’s mechanism design, including the versatile pre-training and the EM-CoT-augmented Fine-tuning. The results are shown in Table 2.
Effectiveness of the versatile pre-training
. As depicted in Panel A, the full versatile pre-training utilizing all three diverse datasets demonstrates unparalleled effectiveness, consistently achieving the highest performance metrics. Conversely, even the selective removal of a single data modality, such as visual generation data (w/o V), leads to a significant degradation in performance, particularly on hard tasks (over 50% reduction). Further progressive exclusion of textual (w/o V+T) and action prediction data (w/o V+T+A) results in a continuously worsening performance trajectory. Notably, without any pre-training (w/o V+T+A), the model’s performance falls to a complete 0% on hard tasks, demonstrating that pre-training is an absolutely foundational requirement for establishing the core competencies necessary to tackle complex problems. Collectively, these findings underscore the indispensable and complementary roles of the versatile pre-training step.
Effectiveness of EM-CoT
. As shown in Panel B, both visual subgoal images and textual reasoning are crucial components for EM-CoT, removing either component degrades performance. Combining both reasoning modalities in Halo achieves the highest accuracy, particularly on hard tasks. These results validate the effectiveness of the proposed EM-CoT formulation.
4.4 Qualitative Results of EM-CoT
To qualitatively assess the efficacy of EM-CoT, we visualize representative rollouts from both the clean and randomized settings of RoboTwin 2.0, as illustrated in Figure 6. In the clean setting, EM-CoT demonstrates high fidelity, with predicted textual reasoning and visual subgoals closely aligning with the ground truth. The result indicates that Halo effectively masters the underlying task logic, providing a clear, self-generated multimodal roadmap that anchors the final action in a semantically and visually consistent manner. Under aggressive environmental randomization, EM-CoT still exhibits significant robustness while maintaining performance despite substantial visual shifts. As shown in the last row of Figure 6, Halo correctly identifies targets and plans trajectories even when object appearances and backgrounds significantly differ from the training distribution. This suggests our EM-CoT paradigm allows the agent to perform true semantic reasoning rather than relying on simple pattern matching. By enforcing a deliberate reasoning process, the model effectively performs low-level control guided by high-level reasoning under the challenging out-of-distribution scenarios.
4.5 Real-World Results
Figure 5: Qualitative Analysis of the EM-CoT. We highlight (i) accurate textual reasoning and subgoal image generation in the clean setting, and (ii) robust EM-CoT generalization under the unseen domain-randomized settings.
Figure 6: Real-World Task Settings. (a) Basic Setting: Four tasks, including tool-use sweeping, bimanual cup nesting, inter-arm screwdriver handover, and placing an object into a drawer. (b) Generalization Setting: Challenging scenarios involving visual distractions (e.g., Coke bottles and cups), lighting variations, background variations (e.g., changing tablecloth colors), and novel objects (e.g., replacing the broom with a sponge).
(a) Basic setting.
(b) Generalization setting.
Figure 7: Quantitative results of real-world experiments on both the basic setting and the generalization setting. Each task is evaluated over 50 trials, and the success rate is reported.
We further evaluate Halo on a real-world Cobot Mobile ALOHA platform on four long-horizon manipulation tasks, including tool-use sweeping, bimanual cup nesting, inter-arm screwdriver handover, and placing an object into a drawer (Figure 6). These tasks require multi-step planning, semantic grounding, and coordinated dual-arm control skills, making them valuable for evaluating the effectiveness of EM-CoT reasoning in real-world settings.
As shown in Figure 7, Halo consistently outperforms the baseline policies $\pi_{0}$ and $\pi_{0.5}$ under both the basic and generalization settings. While the baselines suffer noticeable performance degradation in the presence of visual distractions, lighting and background variations, and novel objects, Halo remains robust and achieves the highest success rates across all tasks. These results demonstrate that EM-CoT reasoning enables more stable long-horizon execution and stronger generalization in real-world robotic manipulation.
5 Conclusion
In this paper, we propose Halo, a unified VLA model that enables embodied multimodal chain-of-thought (EM-CoT) reasoning through joint textual reasoning, visual foresight, and action prediction using a Mixture-of-Transformers (MoT) architecture. We also introduce an automated pipeline for synthesizing EM-CoT training data, along with a carefully crafted training recipe. Extensive experiments and ablation studies validate the effectiveness of Halo in the challenging out-of-distribution and real-world settings, highlighting Halo as a scalable and generalizable human-like VLA reasoning paradigm.
Appendix A Details on Model Architecture
A.1 LLM Backbone
Halo leverages Qwen-2.5-1.5B LLM (Yang et al., 2025) as initial backbone, as illustrated in Table 3.
Table 3: Qwen2.5-1.5B-Instruct Model Configuration Parameters
Parameter Value Parameter Value
architectures [”Qwen2ForCausalLM”] num_attention_heads 12
attention_dropout 0.0 num_hidden_layers 28
bos_token_id 151643 num_key_value_heads 2
eos_token_id 151645 rms_norm_eps $1\times 10^{-6}$
hidden_act ”silu” rope_scaling null
hidden_size 1536 rope_theta 1000000.0
initializer_range 0.02 sliding_window 32768
intermediate_size 8960 tie_word_embeddings true
max_position_embeddings 32768 torch_dtype float32
max_window_layers 21 transformers_version 4.49.0
model_type ”qwen2” use_cache true
vocab_size 151933 use_sliding_window false
A.2 Multimodal Encoders
To jointly train all three modalities, we implement different encoders to map multimodal inputs into tokens, stated as follows:
  • Textual Input: We utilize Qwen2.5’s original tokenizer (Bai et al., 2023).
  • Visual Input: We employ separate visual encoders to handle visual understanding and generation independently. For visual understanding, we utilize a ViT (Dosovitskiy, 2020) encoder initialized with SigLIP2-so400m/14 (Tschannen et al., 2025) (pre-trained at 384×384). To enhance high-resolution perception, we interpolate position embeddings to support inputs up to 980×980 and integrate NaViT (Dehghani et al., 2023) to process images in their native aspect ratios. A two-layer MLP projector is adopted to align visual features with LLM’s embedding space. For visual generation, we leverage the pre-trained VAE (Kingma & Welling, 2013) from FLUX (Labs et al., 2025), featuring a downsample ratio of 8 and a latent channel of 16. The VAE latents are processed by flattening 2×2 spatial patches into vectors, which are then projected by a linear layer to match the hidden dimension of the LLM. Conversely, a VAE decoder is employed similarly to reconstruct images from latents. All VAE parameters remain frozen during training.
  • Action Input: We employ a simple yet effective linear projection layer as the action encoder, mapping the raw continuous actions directly into the LLM’s hidden dimension. Symmetrically, a linear decoder is utilized to project the LLM’s output hidden states back to the original action dimensions for execution.
A.3 Special Tokens
To distinguish different phases of workflow, we add several special tokens to LLM’s vocabulary, listed as follows:
  • $\langle\text{subtask\_start}\rangle$, $\langle\text{subtask\_end}\rangle$
  • $\langle\text{plan\_start}\rangle$, $\langle\text{plan\_end}\rangle$
  • $\langle\text{move\_start}\rangle$, $\langle\text{move\_end}\rangle$
  • $\langle\text{action\_start}\rangle$, $\langle\text{action\_end}\rangle$
  • $\langle\text{vision\_start}\rangle$, $\langle\text{vision\_end}\rangle$
A.4 Attention Mask
In Figure 4, we have elaborated on the attention mask during fine-tuning phase. The attention masking strategy during pre-training phase is illustrated in Figure 8.
Appendix B Details on EM-CoT Data Pipeline
B.1 Action Primitives Extraction
Input: Trajectory data $\mathcal{D}$ with gripper states $G^{l},G^{r}$ and end-effector poses $P^{l},P^{r}$; config thresholds $\theta$
Output: Frame-wise action labels $\mathcal{L}$ with natural language descriptions
// Compute instantaneous idle flags $I^{a}[t]$ for arms $a\in\{l,r\}$:
for $t=1$ to T1T-1 do
 $I^{a}[t]\leftarrow\text{IsIdle}(P^{a}[t],P^{a}[t-1],G^{a}[t],G^{a}[t-1];\theta_{\text{vel}},\theta_{\text{dg}})$;
$I^{a}[0]\leftarrow\text{True}$;
// Extract action segments via long idle periods:
for arm $a\in\{l,r\}$ do
 $S^{a}\leftarrow\text{SegmentActions}(I^{a};\theta_{\text{min\_idle}})$;
// Initialize all frames as idle:
for $t=0$ to T1T-1 do
 $\mathcal{L}[t]\leftarrow\{a:[\text{``idle''},\text{``''}]\mid a\in\{l,r\}\}$;
// Label action segments per arm:
for arm $a\in\{l,r\}$ do
 for each segment $(s,e)\in S^{a}$ do
    // Mark gripper actions:
    for $t=\max(1,s)$ to $e$ do
       if $\Delta G^{a}[t]\leq-\theta_{\text{dg}}$ then
          $\mathcal{L}[t].a\leftarrow[\text{``grasp''},\text{``''}]$;
       else if $\Delta G^{a}[t]\geq\theta_{\text{dg}}$ then
          $\mathcal{L}[t].a\leftarrow[\text{``release''},\text{``''}]$;
    // Label move subsegments:
    for each maximal contiguous idle subsegment $(s^{\prime},e^{\prime})\subseteq(s,e)$ do
       $\Delta P=\sum_{t=s^{\prime}+1}^{e^{\prime}}(P^{a}[t]-P^{a}[t-1])$ ;
       $d\leftarrow\text{GetDirection}(\Delta P;\theta_{\text{dir}})$ ;
       for $t=s^{\prime}$ to $e^{\prime}$ do
          $\mathcal{L}[t].a\leftarrow[\text{``move''},d]$ ;
// Convert labels to natural language:
for $t=0$ to T1T-1 do
 for arm $a\in\{l,r\}$ do
    $\mathcal{L}[t].\text{desc}^{a}\leftarrow\text{LookupSentence}(\mathcal{L}[t].a)$;
return $\mathcal{L}$;
Algorithm 1 Action Primitives Extraction Algorithm
Key Function:
  • IsIdle(): Checks if displacement speed $\theta_{vel}$ and gripper change $\theta_{dg}$.
  • SegmentActions(): Finds motion segments separated by idle periods of $\geq\theta_{min\_idle}$ frames.
  • GetDirection(): Computes dominant motion axes from displacement vector using ratio threshold $\theta_{dir}$.
B.2 Prompts for EM-CoT
Our annotation pipeline employs a three-stage prompting strategy to transform observations into structured, high-level task descriptions. First, a VLM generates a coherent, temporally-ordered narrative of the entire task. Second, this narrative is decomposed into a sequence of goal-oriented subtasks. Finally, each video frame is aligned to its corresponding subtask, with the model providing first-person reasoning that links the observed actions to the overall plan. This staged approach ensures temporal coherence and semantic clarity in the final annotations.
Stage 1: Task Narrative Generation You are an expert roboticist analyzing a human or robot demonstration. Your task is to write a SINGLE, COHERENT, HIGH-LEVEL NARRATIVE paragraph that STRICTLY follows the TEMPORAL ORDER. Overall Goal: <instruction> Per-frame low-level actions are provided for context only—DO NOT copy or list them. Low-level Arm Actions: 1. Frame_id:0, Left arm action:<left_description_0>, Right arm action:<right_description_0> 2. Frame_id:1, Left arm action:<left_description_1>, Right arm action:<right_description_1> 3. $\dots$ Instructions: 1. Describe the task step by step in exact chronological order. 2. Explicitly state simultaneous bimanual actions (e.g., “At the same time, the left hand stabilizes… while the right hand unscrews…”). 3. Focus on purpose: explain what each action achieves toward the goal. 4. Use specific object names when identifiable. Output Rules: Produce exactly one fluent paragraph (2–4 sentences). Output ONLY the narrative. No markdown, bullets, or extra text.
Stage 2: Subtask Sequence Extraction You are an expert in robotic task analysis. Based on the task narrative below, decompose the task into a sequence of HIGH-SEMANTIC, GOAL-ORIENTED SUBTASKS. Task Narrative: <narrative> Instructions: 1. Split the narrative into discrete subtasks in strict chronological order. 2. Represent coordinated bimanual actions as ONE subtask (never split). 3. Express high-level intent (e.g., “Assemble the lid onto the container”), not low-level motions (e.g., “move”, “grab”). 4. Use imperative, active voice; keep the list short (typically 2–5 subtasks). Output Format: Return ONLY a JSON list of strings. Example: ["Pick up red cup", "Pour water into cup"]
Stage 3: Subtask Alignment and Reasoning You are the autonomous onboard controller of a robot. You are currently executing a task. Describe your reasoning in the FIRST PERSON (“I”, “me”). Overall Goal: <instruction> Planned Subtask Sequence: 1. <subtask_1> 2. <subtask_2> 3. $\dots$ Low-level Arm Actions: 1. Frame_id:0, Left arm action:<left_description_0>, Right arm action:<right_description_0> 2. Frame_id:1, Left arm action:<left_description_1>, Right arm action:<right_description_1> 3. $\dots$ Instructions: 1. For each segment, explain your internal decision-making logic from a first-person view. 2. Include: (a) visual observation, (b) goal-driven inference, (c) movement logic. 3. Ensure physical alignment: do not claim a subtask has started if low-level logs show idle. 4. Keep reasoning under 50 words per segment; account for every frame. Output Format: Return ONLY a JSON list of objects with keys: "subtask", "frame" (as [start, end]), and "reasoning".
B.3 Visual Subgoal Extraction
Input: Frame annotations $\mathcal{A}=\{a_{0},a_{1},\dots,a_{T-1}\}$, where $a_{t}=(\texttt{frame\_id}=t,\texttt{subtask}=s_{t})$;
Image sequence $\mathcal{I}=[I_{0},I_{1},\dots,I_{T-1}]$
Output: Goal image sequence $\mathcal{G}=[G_{0},G_{1},\dots,G_{T-1}]$
// Step 1: Identify goal frame for each subtask
Initialize map $\mathcal{M}\leftarrow\emptyset$;
, Sort $\mathcal{A}$ by frame_id ;
for $t\leftarrow 1$ to T1T-1 do
 if $s_{t}\neq s_{t-1}$ then
    $\mathcal{M}[s_{t-1}]\leftarrow t$ ;
$\mathcal{M}[s_{T-1}]\leftarrow T-1$ ;
// Step 2: Broadcast goal images to all frames
for $t\leftarrow 0$ to T1T-1 do
 $s\leftarrow s_{t}$ ;
   , $t_{g}\leftarrow\mathcal{M}[s]$ ;
   , $G_{t}\leftarrow I_{t_{g}}$ ;
return $\mathcal{G}$;
Algorithm 2 Visual Subgoal Extraction from Subtask Boundaries
Appendix C Training Implementation
In this section, we provide details on pre-training and fine-tuning of Halo. Halo is pre-trained on 32 Nvidia H100 GPUs for 90k steps, with a sequence length of 40k per rank. Then, we fine-tune Halo on 32 Nvidia H100 GPUs with sequence length of 27k, 110k steps for simulation, 80k steps for real experiment. We illustrate the detailed hyperparameters in Table 4.
Table 4: Training Hyperparameters. Detailed configurations for the two-stage training process of Halo.
Configuration Pre-training Fine-tuning
Base Architecture Qwen2.5-1.5B $\times$ 3 Experts
Total Parameters $\approx$ 4.5B
Optimizer AdamW
Learning Rate $1\times 10^{-4}$ $5\times 10^{-5}$
Learning Rate Schedule Constant Constant
Weight Decay 0.0 0.0
Per rank Sequence Length 40k 27k
Warm-up Steps 2000 500
Training Steps 90k 110k / 80k
Und Resolution (378, 980)
Gen Resolution (512, 1024)
Loss Weight CE:MSE:L1 = 1:2:4 CE:MSE:L1 = 1:1:1
Appendix D Complete Result on RoboTwin2.0 Benchmark
See Table 5 for complete result on RoboTwin 2.0 benchmark.
Table 5: Complete Result on RoboTwin
Task $\pi_{0}$ RDT-1B Diffusion Policy Halo-w/o CoT Halo-w/ CoT
Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard
Adjust Bottle 90 56 81 75 97 0 99 2 97 9
Beat Block Hammer 43 21 77 37 42 0 95 3 96 11
Blocks Ranking RGB 19 5 3 0 0 0 85 3 94 7
Blocks Ranking Size 7 1 0 0 1 0 37 0 58 2
Click Alarmclock 63 11 61 12 61 5 85 10 83 14
Click Bell 44 3 80 9 54 0 100 5 100 10
Dump Bin Bigbin 83 24 64 32 49 0 88 30 93 28
Grab Roller 96 80 74 43 98 0 97 49 95 57
Handover Block 45 8 45 14 10 0 84 24 81 36
Handover Mic 98 13 90 31 53 0 94 61 96 61
Hanging Mug 11 3 23 16 8 0 33 2 48 5
Lift Pot 84 36 72 9 39 0 94 23 95 34
Move Can Pot 58 21 25 12 39 0 90 11 94 15
Move Pillbottle Pad 21 1 8 0 1 0 69 10 76 26
Move Playingcard Away 53 22 43 11 47 0 91 40 89 53
Move Stapler Pad 0 2 2 0 1 0 47 11 45 19
Open Laptop 85 46 59 32 49 0 81 23 89 37
Open Microwave 80 50 37 20 5 0 71 5 86 24
Pick Diverse Bottles 27 6 2 0 6 0 59 14 76 17
Pick Dual Bottles 57 12 42 13 24 0 75 28 82 30
Place A2B Left 31 1 3 1 2 0 55 8 68 8
Place A2B Right 27 6 1 1 13 0 54 7 52 9
Place Bread Basket 17 4 10 2 14 0 85 14 90 26
Place Bread Skillet 23 1 5 1 11 0 78 19 85 23
Place Burger Fries 80 4 50 27 72 0 98 33 99 37
Place Can Basket 41 5 19 6 18 0 60 33 68 34
Place Cans Plasticbox 34 2 6 5 40 0 99 40 98 47
Place Container Plate 88 45 78 17 41 0 93 18 96 22
Place Dual Shoes 15 0 4 4 8 0 17 2 15 3
Place Empty Cup 37 11 56 7 37 0 92 24 95 28
Place Fan 20 10 12 2 3 0 47 6 62 9
Place Mouse Pad 7 1 1 0 0 0 53 8 51 12
Place Object Basket 16 2 33 17 15 0 82 20 89 25
Place Object Scale 10 0 1 0 1 0 41 2 55 5
Place Object Stand 36 11 15 5 22 0 79 29 84 33
Place Phone Stand 35 7 15 6 13 0 83 6 91 10
Place Shoe 28 6 35 7 23 0 72 13 70 18
Press Stapler 62 29 41 24 6 0 86 58 92 64
Put Bottles Dustbin 54 13 21 4 22 0 73 8 80 13
Put Object Cabinet 68 18 33 18 42 0 44 4 59 8
Rotate QRcode 68 15 50 5 13 0 63 8 69 11
Scan Object 18 1 4 1 9 0 64 19 73 24
Shake Bottle Horizontally 99 51 84 51 59 18 100 63 100 66
Shake Bottle 97 60 74 45 65 8 100 70 98 73
Stack Blocks Three 17 0 2 0 0 0 90 32 96 37
Stack Blocks Two 42 1 21 2 7 0 99 55 100 60
Stack Bowls Three 66 24 51 17 63 0 83 20 92 25
Stack Bowls Two 91 41 76 30 61 0 96 46 98 49
Stamp Seal 3 4 1 0 2 0 49 16 60 21
Turn Switch 27 23 35 15 36 1 58 24 65 27
Average 46.4 16.3 34.5 13.7 28.0 0.6 75.3 21.2 80.5 26.4
Appendix E Details on Real-World Experiment
Our real-world experiments are all conducted with COBOT Magic, a robot using the Mobile ALOHA system design (Fu et al., 2024) and manufactured by AgileX Robotics. Notably, we exclusively utilize the robot’s head camera as the sole source of visual input, requiring the model to reason and act based on an egocentric perspective. To evaluate policies, we design a set of tasks that span various dimensions, including:
  • Tool-mediated Manipulation: Gently pick up the broom and sweep the buttons on the board into the dustpan. (”Sweep the buttons”)
  • Semantic Visual Grounding: Place the grey cup into the pink cup with left gripper, and the green cup into the yellow cup with right gripper. (”Bimanual cup nesting”)
  • Inter-arm Collaboration: Handover the screwdriver from the left to the right. (”Handover the screwdriver”)
  • Multi-step Manipulation: Open the top drawer, put the lemon on the plate into the top drawer, close the drawer. (”Put lemon into drawer”)
We collect 80 trajectories for each task, and process them through our automated pipeline to yield the EM-CoT data. We then aggregate the data across all four tasks into a multi-task corpus for finetuning.
In order to evaluate generalization, we design settings that span various axes of unseen scenarios, including visual distraction, lightning variation, background variation, and novel object. The generalization settings are formulated as follows.
  • Visual Distraction: Randomly placing irrelevant objects such as Coke bottle, cups onto the table.
  • Lightning Variation: Varying the illumination levels of a desk lamp to introduce visual interference.
  • Background Variation: Changing the color of the tablecloth.
  • Novel Object: For the task of sweeping the buttons, we change the broom to a sponge during evaluation.
Figure 8: Attention map during pre-training phase.